# Cascade of topological phase transitions and revival of topological zero modes in imperfect double helical liquids

### v1

@@ VI Discussion
From an experimental perspective, we demonstrate that random backscattering is not merely a detrimental perturbation; it also provides an additional tuning knob that can selectively activate one branch of the Majorana zero mode criterion. Moreover, we demonstrate that at moderately strong interactions, pairing asymmetry suppresses the $N_{\mathrm{mzm}}=2$ phase, leading to additional topological phases and resulting in a cascade of transitions and a revival of a Majorana zero mode. These features can be electrically accessible by controlling the screened Coulomb interactions in the edge channels. This suggests that systems with spin-flip backscattering, such as helical channels in quantum spin Hall devices with dilute magnetic impurities or in-plane fields combined with charge disorder, can serve as an equally suitable and potentially more tunable platform for topological electronics [45].
@@
Beyond transport signatures, direct detection of the Majorana zero modes stabilized in our setting can be achieved through local spectroscopic probes. Scanning tunneling spectroscopy with high energy resolution, as employed to resolve zero-bias bound states in iron-based superconductors [132], provides an experimentally accessible route to identifying the zero modes predicted here. The evolution of the local density of states under controlled tuning of in-plane fields, interaction strength, or channel asymmetries offers a concrete path toward verifying the disorder-enabled topological windows, the cascades of transitions, and the revival of Majorana zero modes in future device platforms.
@@ Acknowledgements.
We thank H.-C. Wang for interesting discussions. This work was financially supported by the National Science and Technology Council (NSTC), Taiwan, through Grant No. NSTC-112-2112-M-001-025-MY3 and Grant No. NSTC-114-2112-M-001-057, and Academia Sinica (AS), Taiwan through Grant No. AS-iMATE-114-12. C.-H.H. acknowledges support from the National Center for Theoretical Sciences (NCTS), Taiwan.
@@ Appendix A Potential platforms
In this section, we compile material parameters of several established and newly emerging quantum spin Hall systems in Tables 13. These include semiconductor-based materials, monolayers, and twisted bilayers that host (fractional) quantum spin Hall states. Conventional semiconductor systems have small to moderate bulk gaps and support quantized edge transport at sufficiently low temperatures. Monolayer systems, such as WTe2, TaIrTe4 and bismuthene on SiC, typically exhibit larger gaps. In particular, TaIrTe4 is notable for hosting a density-tunable dual quantum spin Hall insulating phase [144], with quantized edge conduction observed up to nearly 100 K. These large-gap and high-temperature systems offer promising conditions for robust helical edge states.
@@
[…]
@@
Bulk gap, $\Delta=\hbar v_{F}/a$ 44 meV ($\nu=2$) $0.3$ meV ($\nu=3$)
@@ Appendix C Derivation of RG flow equations
$\quad\times\left(1-\frac{|z|^{2}}{a^{2}}a^{2}(\lambda_{\phi}^{2}J_{\phi}\bar{J}_{\phi}-\lambda_{\theta}^{2}J_{\theta}\bar{J}_{\theta})+\dots\right).$
@@
with $\mathbf{r}_{1,2}\equiv(r_{1,2},y_{1,2})$, $y=u_{n}\tau$ and the partition function $\mathcal{Z}_{n}\equiv\int\mathcal{D}\phi_{n}e^{-(S_{{\rm el},n}+S_{{\rm rs},n})/\hbar}$, and under the rescaling the correlator preserves the helical liquid form with renormalized parameters.
@@ D.2 Majorana zero modes and topological criterion
$\Phi_{j}^{<}(r)$ $\equiv$ $\Psi_{j}^{<}e^{\kappa^{<}r},\qquad r<0,$ (61a)
$\Phi_{j}^{>}(r)$ $\equiv$ $\Psi_{j}^{>}e^{-\kappa^{>}r},\qquad r>0.$ (61b)
@@
up to normalization prefactors. The lower Nambu components follow from particle-hole self-conjugation, as enforced in Eq. (59), and are therefore not shown.
Figure 11: Topological phase diagram constructed based on $N_{\text{mzm}}$ as a function of $\Delta_{c}/\Delta_{+}$ and $B_{+}/\Delta_{+}$. Colors indicate $N_{\text{mzm}}$ computed from Eq. (69); regions with $N_{\text{mzm}}=1$ are further resolved according to which of the $\varepsilon$ conditions in Eq. (70) is fulfilled.
@@
With the above formulation, we numerically evaluate $N_{\mathrm{mzm}}$ across parameter space to construct the topological phase diagram within the single-particle description. A series of examples is given in Fig. 11. The phase boundaries coincide precisely with the bulk gap-closing conditions in Eq. (56), with straight lines arising from the decoupled region ($\bar{\Delta}_{c}=0$) and curved lines from the coupled region ($\bar{\Delta}_{c}\neq 0$).

### v2

@@ VIII Discussion
From an experimental perspective, we demonstrate that random backscattering is not merely a detrimental perturbation; it also provides an additional tuning knob that can selectively activate one branch of the Majorana zero mode criterion. Moreover, we demonstrate that at moderately strong interactions, pairing asymmetry suppresses the $N_{\mathrm{mzm}}=2$ phase, leading to additional topological phases and resulting in a cascade of transitions and a revival of a Majorana zero mode. These features can be electrically accessible by controlling the screened Coulomb interactions in the edge channels. Our analysis is also relevant to phase-controlled Josephson junctions, where access to topological phases relies on phase biasing between two channels via an external magnetic flux [139, 140, 141, 142]. In such devices, imperfect magnetic shielding can in practice introduce residual fields in the conduction channels, effectively generating spin-flip backscattering. Taken together, these considerations suggest that systems with controlled spin-flip backscattering, such as helical channels in quantum spin Hall devices with dilute magnetic impurities or in-plane fields combined with charge disorder, can serve as a suitable and potentially more tunable platform for topological electronics [50].
@@
Beyond transport signatures, the Majorana zero modes stabilized in our setting can be directly probed using local probes. Scanning tunneling spectroscopy with high energy resolution, as demonstrated in the detection of zero-bias bound states in iron-based superconductors [135], provides an experimentally accessible route to identifying the zero modes predicted here. As revealed by the density-profile analysis, the spatial localization of the zero modes constitutes an alternative scanning-probe feature, with pronounced broadening indicating proximity to a topological phase transition. Overall, the evolution of the local density of states and spatial density profiles under controlled tuning of in-plane fields, interaction strength, or channel asymmetry offers a concrete pathway to experimentally verify the disorder-enabled topological windows, cascades of phase transitions, and the revival of Majorana zero modes in future device platforms.
@@ Acknowledgements.
We thank A. Alexandradinata, B. A. Bernevig, Y.-Y. Chang, A. Garg, T. Grover, H.-H. Lu, D. Parker, S. Ryu, S. Syzranov, H.-C. Wang, and T. Yoshida for interesting discussions. This work was financially supported by the National Science and Technology Council (NSTC), Taiwan, through Grant No. NSTC-112-2112-M-001-025-MY3 and Grant No. NSTC-114-2112-M-001-057, and Academia Sinica (AS), Taiwan through Grant No. AS-iMATE-114-12.
Data Availability
The data that support the findings of this study are available at Zenodo [145].
@@ Appendix A Potential platforms
In this section, we compile material parameters of several established quantum spin Hall materials and newly emerging quantum spin Hall systems. These include semiconductor-based materials (see Table 1), monolayers (see Table 2), and twisted bilayers that host (fractional) quantum spin Hall states (see Table 3). Conventional semiconductor systems typically have small to moderate bulk gaps and support quantized edge transport at sufficiently low temperatures for small sample size. Monolayer systems, such as WTe2, TaIrTe4 and bismuthene on SiC, typically exhibit larger gaps. In particular, TaIrTe4 is notable for hosting a density-tunable dual quantum spin Hall insulating phase [157], with quantized edge conduction observed up to nearly 100 K. These large-gap and high-temperature systems offer promising conditions for robust helical edge states.
@@
Edge-state Fermi velocity, $v_{F}$
$1.2\times 10^{5}$~m/s (Y-edge)
@@
Bulk gap, $\Delta=\hbar v_{F}/a$
44 meV ($\nu=2$)
@@ Appendix C Derivation of RG flow equations
$\quad\times\left(1-\frac{|z|^{2}}{a^{2}}a^{2}(\lambda_{\phi}^{2}J_{\phi}\bar{J}_{\phi}+\lambda_{\theta}^{2}J_{\theta}\bar{J}_{\theta})+\dots\right).$
@@
with $\mathbf{r}_{1,2}\equiv(r_{1,2},y_{1,2})$, $y=u_{n}\tau$ and the partition function $\mathcal{Z}_{n}\equiv\int\mathcal{D}\phi_{n}e^{-(S_{{\rm el},n}+S_{{\rm rs},n})/\hbar}$. Under the rescaling, the correlator preserves the helical liquid form with renormalized parameters. In the above, the pairing terms are omitted in $\mathcal{Z}_{n}$ as the crossed-term contributions to the RG flow vanish; see Appendix C.2.
@@ C.1 RG procedure for general m
C.2 Contributions from the crossed terms
We now consider the second-order expansion of the partition function,
$\mathcal{Z}=\mathcal{Z}_{0}\left[1-\left\langle\frac{\delta S}{\hbar}\right\rangle_{0}+\frac{1}{2}\left\langle\left(\frac{\delta S}{\hbar}\right)^{2}\right\rangle_{0}+\dots\right],$ (54)
where $\delta S=S_{s,1}+S_{s,2}+S_{\times}+\dots$, corresponding to the non-quadratic terms in the effective action. The contributions from the non-crossed terms, such as $\langle S_{\mathrm{s},1}^{2}/\hbar^{2}\rangle_{0}$, have been discussed above. Here we explicitly examine the crossed terms between the local and nonlocal pairings,
$\frac{S_{\mathrm{s},1}}{\hbar}=\frac{\tilde{\Delta}_{1}}{\pi a^{2}}\int d^{2}\mathbf{x}\,\cos(2\sqrt{m}\tilde{\theta}_{1}),$ (55)
$\frac{S_{\times}}{\hbar}=\frac{\tilde{\Delta}_{c}}{\pi a^{2}}\int d^{2}\mathbf{x}\left\{\cos\left[\sqrt{m}(-\tilde{\phi}_{1}+\tilde{\theta}_{1}+\tilde{\phi}_{2}+\tilde{\theta}_{2})\right]\right$.
$\qquad\qquad\qquad\qquad\left.+\cos\left[\sqrt{m}(\tilde{\phi}_{1}+\tilde{\theta}_{1}-\tilde{\phi}_{2}+\tilde{\theta}_{2})\right]\right\}.$
We first evaluate the OPE for the contraction of $S_{\mathrm{s},1}$ with the first term of $S_{\times}$. In contrast to Eq. (37), the field mismatch in the crossed terms yields a global prefactor,
$e^{i2\sqrt{m}\tilde{\theta}_{1}(z_{1},\bar{z}_{1})}e^{-i\sqrt{m}(-\tilde{\phi}_{1}(z_{2},\bar{z}_{2})+\tilde{\theta}_{1}(z_{2},\bar{z}_{2})+\tilde{\phi}_{2}(z_{2},\bar{z}_{2})+\tilde{\theta}_{2}(z_{2},\bar{z}_{2}))}$ (56)
$+e^{-i2\sqrt{m}\tilde{\theta}_{1}(z_{1},\bar{z}_{1})}e^{i\sqrt{m}(-\tilde{\phi}_{1}(z_{2},\bar{z}_{2})+\tilde{\theta}_{1}(z_{2},\bar{z}_{2})+\tilde{\phi}_{2}(z_{2},\bar{z}_{2})+\tilde{\theta}_{2}(z_{2},\bar{z}_{2}))}$
$\approx\frac{2}{\left(\frac{L}{a}\right)^{\frac{m}{4}\left(K_{1}+\frac{1}{K_{1}}+K_{2}+\frac{1}{K_{2}}\right)}\left(\frac{|z|}{a}\right)^{\frac{m}{K_{1}}}}$
$\quad\times\left(1-\frac{|z|^{2}}{a^{2}}a^{2}(\frac{9m}{4}J_{\phi}\bar{J}_{\phi}+\frac{m}{4}J_{\theta}\bar{J}_{\theta})+\dots\right),$
dependent on the system size $L$. To find the contribution to the RG flow of $K_{1}$, we integrate the relative coordinate over the thin shell $a\leq|z|\leq a(1+dl)$, leading to
$\frac{dK_{1}}{dl}\propto\left(\frac{L}{a}\right)^{-\frac{m}{4}(K_{1}+\frac{1}{K_{1}}+K_{2}+\frac{1}{K_{2}})}\tilde{\Delta}_{1}\tilde{\Delta}_{c}.$ (57)
Since the scaling exponent is strictly positive, this RG contribution is strongly suppressed for sufficiently large $L/a$. Similarly, the other crossed term combinations are also suppressed and the RG flow of $K_{n}$ is dominated by the non-crossed-term contributions presented in Eq. (11d). Similarly, crossed-term corrections to the couplings, such as $d\tilde{\Delta}_{c}/dl\propto\tilde{\Delta}_{1}\tilde{\Delta}_{c}$) are also strongly suppressed.
@@ D.2 Majorana zero modes and topological criterion
$\Phi_{j}^{<}(r)$ $\equiv$ $\Psi_{j}^{<}e^{\kappa^{<}_{j}r},\qquad r<0,$ (68a)
$\Phi_{j}^{>}(r)$ $\equiv$ $\Psi_{j}^{>}e^{-\kappa^{>}_{j}r},\qquad r>0.$ (68b)
@@
up to normalization prefactors. In the above, the four solutions with $(n,\epsilon)$ for $n\in\{1,2\}$ and $\epsilon\in\{+,-\}$ correspond to the four $\Psi_{j}^{<}$’s for $j\in\{1,\cdots,4\}$ in Eq. (68a). The lower Nambu components follow from particle-hole self-conjugation, as enforced in Eq. (66), and are therefore not shown.
Figure 14: Topological phase diagram constructed based on $N_{\text{mzm}}$ as a function of $\Delta_{c}/\Delta_{+}$ and $B_{+}/\Delta_{+}$. Colors indicate $N_{\text{mzm}}$ as computed from Eq. (76) and verified numerically; regions with $N_{\text{mzm}}=1$ are further resolved according to which of the $\varepsilon$ conditions in Eq. (77) is fulfilled. The solid (curved) and dashed (horizontal) lines coincide with the analytical phase boundaries given in Eq. (63) of the coupled ($\bar{\Delta}_{c}\neq 0$) and decoupled ($\bar{\Delta}_{c}=0$) regions, respectively.
@@
In the above, the four solutions with $(\lambda,\epsilon)$ for $\lambda,\epsilon\in\{+,-\}$ correspond to the four $\Psi_{j}^{>}$’s for $j\in\{1,\cdots,4\}$ in Eq. (68b).
@@
With the above formulation, we numerically evaluate $N_{\mathrm{mzm}}$ across parameter space to construct the topological phase diagram within the single-particle description. A series of examples is given in Fig. 14. The phase boundaries coincide precisely with the bulk gap-closing conditions in Eq. (63), with straight lines arising from the decoupled region ($\bar{\Delta}_{c}=0$) and curved lines from the coupled region ($\bar{\Delta}_{c}\neq 0$). Thus, the full phase diagram is determined by the superposition of the phase boundaries from both regions. The topological invariant $N_{\text{mzm}}$ can change only when one of these boundaries is crossed.
@@ E.5 Numerical results for fractional helical liquids
Figure 23: Channel-resolved density profiles $\rho^{(n)}(r)\equiv\sum_{j}|\Psi_{\text{mzm},j}^{(n)}(r)|^{2}$ of the Majorana zero modes along the local coordinate $r$. From (a) to (c), we have $K_{+}(0)=0.54,\,0.57$ [identical to $P$ in Fig. 7(b)] and $0.61$, respectively. The red (orange) curves show the density for the channel $n=1$ ($n=2$). The other parameters include $\tilde{\Delta}_{+}(0)=0.03$, $\tilde{\Delta}_{-}(0)=0.01$, $\tilde{\Delta}_{c}(0)=0.01$, $\tilde{D}_{+}(0)=10^{-6}$ and $\tilde{D}_{-}(0)=K_{-}(0)=0$; see Table 4 for the complete sets of the adopted parameter values.
E.6 Representative spatial density profiles of the zero modes
To demonstrate the spatial density profile of the decaying modes, we show the numerical results for three representative values of $K_{+}(0)$ in Fig. 23 and plot the channel resolved density profiles. Figs. 23(a–c) show the resulting channel-resolved density profiles for the corresponding values of $K_{+}(0)$. Figs. 23(a,c) correspond to the $N_{\rm mzm}=2$ regions, while Fig. 23(b) is in the intermediate region where one of the modes vanishes, $N_{\rm mzm}=1$. In the resulting plots, we observe exponentially decaying amplitudes. Consistent with the analysis in the main text, we note sharper localization on the $r>0$ side and slower decay on the other side. We can also observe broadening of the profiles as the interactions get weaker.